# Implications of portal vector-like lepton on associated Higgs production at a multi-TeV muon collider

### v1

@@ Abstract
Furthermore, this process can also be used to constrain the dark photon mass ($m_{\gamma_{d}}$) and/or the dark gauge coupling ($g_{d}$) within the pVLL framework. We perform a detailed collider analysis of the $h\gamma_{d}$ process in the $b\bar{b}+$ missing energy final state. A $2\sigma$ exclusion limit for $m_{\gamma_{d}}$ up to $75$ GeV is obtained assuming $g_{d}=0.05$, $\sin\theta_{L}=4\times 10^{-5}$, $\sin\theta_{s}=0.05$, and heavy lepton mass $\gtrsim 2$ TeV at a $3$ TeV muon collider with an integrated luminosity of $1$ ab-1.
@@ 1 Introduction
In recent years, the idea of VLLs as portal matter [42, 43, 44, 18] or portal VLLs (pVLLs) has gained significant interest owing to their rich phenomenology, their ability to evade existing collider constraints, and their potential connection to the dark sector, particularly dark matter phenomenology.
@@
The main highlight of this work is that the $h\gamma_{d}$ process not only yields a higher Higgs production rate than $hZ$ production at a muon collider, but also offers a noble way to explore dark photons in the mass range 1010100100 GeV, independent of the gauge kinetic mixing parameter ($\varepsilon$) [49, 50, 51, 52]. This is because of the fact that the muon–heavy muon–Higgs vertex is sensitive to the dark photon mass ($m_{\gamma_{d}}$) and the dark gauge coupling ($g_{d}$). Consequently, the absence of any observable enhancement in Higgs production along with missing energy at a muon collider would imply a stringent limit on $m_{\gamma_{d}}$ or $g_{d}$, for fixed values of other relevant model parameters. We illustrate this using several representative benchmark points and show that, even for fermion mixing angle as small as $\sim 10^{-6}$ and heavy muon masses in the range 1133 TeV, a wide range of dark photon masses can be excluded.
@@ 2 Theoretical Framework
We consider a framework in which the Standard Mode is extended by a vector-like lepton, singlet under $SU(3)_{C}\times SU(2)_{L}$ with $U(1)_{Y}$ hypercharge 1-1. The VLL plays the role of a portal matter, linking the visible SM sector with a dark sector. A simplified realization of such a scenario is to have a VLL transforming non-trivially under an additional local $U(1)_{D}$ symmetry of the dark sector with corresponding gauge boson referred to as the dark photon ($\gamma_{d}$). The dark photon acquires mass through spontaneous breaking of $U(1)_{D}$, achieved by introducing a complex scalar field $\Phi_{d}$ which is a singlet under the SM gauge group but charged under $U(1)_{D}$. Furthermore, if $\Phi_{d}$ carries the same $U(1)_{D}$ charge as the pVLL, a Yukawa-type portal interaction $\omega_{f}\,\Phi_{d}\,\overline{\mu}^{\prime}_{p_{L}}\mu^{\prime}_{R}$ is permitted. The Yukawa interaction reduces to a bilinear term between the pVLL and the SM lepton once $\Phi_{d}$ gets a non-zero VEV. In this work, we focus on the case where the pVLL mixes with the second-generation charged lepton only to avoid constraints coming from flavor violating observables. The assignment of charge of the relevant field content is presented in Table 1.
@@ 2.3 Electroweak precision constraints
The presence of VLL has various observable consequences and can potentially modify the predictions of the SM. It can contribute to the modification of muon life-time, the branching ratio of $Z\to\mu^{+}\mu^{-}$, total width of the Z and W bosons and correction to the oblique parameters etc. Since the VLL carries electromagnetic charge it can modify the Higgs to di-photon decay rate as well. In our model, the pVLL can couple to the Higgs via muon-heavy muon mixing and/or Higgs-dark Higgs mixing. Therefore, the corresponding interaction vertices will be proportional to $\sim(y_{m}\cos\theta_{R}\cos\theta_{L}\cos\theta_{s}\penalty 10000\ +\penalty 10000\ \omega_{f}\cos\theta_{R}\sin\theta_{L}\sin\theta_{s}$). Hence, for small $\sin\theta_{L}$ and $\sin\theta_{s}$ the deviation from the SM $h-\mu-\mu$ vertex factor is less significant. The correction to the oblique parameters, in particular, S and T due to the singlet VLL is expressed as [71]
@@ 3.1 $hZ$ vs $h\gamma_{d}$ at muon collider
Figure 2: Ratios of cross sections ($R_{i}$’s defined in Eq. (21) for different benchmark points at (a) $\sqrt{s}=3$ TeV and (b) $10$ TeV.
One can see that while the cross section for $\mu^{+}\mu^{-}\to hZ$ production in this scenario remains the same as that in the SM, a higher Higgs production rate is predicted in the $\mu^{+}\mu^{-}\to h\gamma_{d}$ channel. This is true even for a $\sin\theta_{L}$ as small [^4] Small $\sin{\theta_{L}}$ ($\sim 10^{-4}$ or less) is favoured by perturbative unitarity requirements for heavy muon mass beyond TeV scale as $\sim 10^{-5}$ with reasonable choices of other relevant parameters (see Table 2 for specification of benchmark points). The underlying reasons behind this is that $\mu_{R}^{\prime}$ and $\mu_{p_{{}_{R}}}^{\prime}$ share identical $SU(2)_{L}\times U(1)_{Y}$ charges. Hence, after transforming them into mass eigen-basis using Eq. (10) no off-diagonal $\mu^{-}_{R}-\mu^{-}_{p_{{}_{R}}}-Z$ interaction arises. However, the absence of $U(1)_{D}$ charge for the muon induces an off-diagonal $\mu_{R}^{-}-\mu^{-}_{p_{{}_{R}}}-\gamma_{d}$ interaction in the mass basis. It turns out that the corresponding interaction vertex $\mu^{+}_{L}-\mu_{p_{{}_{R}}}^{-}-Z$ is not present in this model while the $\mu^{+}_{L}-\mu_{p_{{}_{R}}}^{-}-\gamma_{d}$ (or $\mu^{-}_{R}-\mu_{p_{{}_{L}}}^{+}-\gamma_{d}$ ) vertex factor is proportional to $\sim g_{d}\sin\theta_{R}\cos\theta_{R}$ (See Eq. A.2). The mixing angle ($\sin\theta_{R}$) between the right chiral muon and heavy muon is given by
@@
In Fig. 3, we plot $R_{2}$ which represent the ratio of average unpolarized cross section for $\mu^{+}\mu^{-}\to h\gamma_{d}$ to that of SM $\mu^{+}\mu^{-}\to hZ$ (defined in Eq. (21)) in the $m_{\gamma_{d}}$$m_{\mu_{p}}$ plane at $\sqrt{s}=3$ TeV to illustrate the effect of these two parameters on $R_{2}$ and hence the $h\gamma_{d}$ cross section for two different $g_{d}$ and $\sin\theta_{L}$ and fixed $\sin\theta_{s}$. The perturbative unitarity limit obtained using Eq. (15) is also displayed on the same plot. Since the requirement of perturbative unitarity disfavors low values of $v_{d}$, one cannot achieve an arbitrary high values of $R_{2}$ by going to lower values of dark photon mass or higher values of $g_{d}$ and a trade off is maintained in the choice of the two independent parameters $m_{\gamma_{d}}$ and $g_{d}$}.
Figure 3: Variation of ratio of cross sections ($R_{2}$) for the process $\mu^{+}\mu^{-}\to h\gamma_{d}$ w.r.t SM $hZ$ production rate in $m_{\gamma_{d}}$$m_{\mu_{{}_{p}}}$ plane assuming $\sqrt{s}=3$ TeV for two different set of ($g_{d},\penalty 10000\ \sin\theta_{L}$) and a fixed value of $\sin\theta_{s}$. (a) The left plot corresponds to $g_{d}=0.05,\penalty 10000\ \sin\theta_{s}=0.05,\penalty 10000\ \text{and}\penalty 10000\ \sin\theta_{L}=4\times 10^{-5}$ and (b) the right plot corresponds to $g_{d}=0.5,\penalty 10000\ \sin\theta_{s}=0.05,\penalty 10000\ \text{and}\penalty 10000\ \sin\theta_{L}=1.66\times 10^{-6}$. The dashed (white) line corresponds to perturbative unitarity constraint on $\omega_{f}$ as described by Eq. (15).
@@ 3.2 Analysis of $h\gamma_{d}$ process
The cross section for $h\gamma_{d}$ production strongly depends on dark photon mass and dark gauge coupling as well as on other model parameters (see Fig. 3 and also the corresponding vertex factors in the Appendix A). The cross section for the $\mu^{+}\mu^{-}\to h\gamma_{d}$ process as a function of the dark gauge coupling ($g_{d}$) can be parametrized as
@@
As already mentioned, Higgs boson produced at the future muon collider center-of-mass energies in a $2\to 2$ process are typically highly boosted. The Higgs subsequently decays into a pair of bottom quarks which due to their small angular separation, manifest as a single collimated jet in the detector. In Fig. 9 and 10, we plot the transverse momentum distributions of a jet or di-jet system, as applicable, for two different muon collider center-of-mass energies. The purpose of these distributions are two-fold. They provide an explanation of Fig 7 and 8 and also give a motivation for the following discussion.
@@ 3.3 Kinematic variables
    In Figure 11, we plot the missing energy distributions for the signal events for various benchmark scenarios corresponding to two different final states at center-of-mass energy $\sqrt{s}=3\penalty 10000\ \text{TeV}$. The corresponding distributions at 10 TeV center-of-mass energy are shown in Fig. 12. The contribution to the missing energy for the signal events is due to the invisible dark photon and the distribution is expected to peak $\sim{\sqrt{\rm s}/2}$ (neglecting the finite mass effect) in a $2\to 2$ process. However, due to the presence of additional neutrinos coming from $B$ meson decays the missing energy distribution extends beyond $\sqrt{\rm s}/2$}. The SM background events in the $b\bar{b}\nu\bar{\nu}$ final state has dominant contribution from the $2\to 3$ processes such as $h\nu\bar{\nu}$ and $Z\nu\bar{\nu}$. Together they form the 99 % of the total cross section and only a very small fraction of it comes from $2\to 2$ processes like $ZZ$ and $hZ$. The pair of neutrinos in the $h\nu\bar{\nu}$ or $Z\nu\bar{\nu}$ processes carries maximum share of the available center-of-mass energy. Hence, the missing energy distribution for the SM background events has a peak mostly around $\sqrt{s}$. This is reflected in the missing energy distributions for individual SM background components and the total SM background as depicted in the same figures. Therefore, missing energy variable is an extremely useful discriminator between the signal and the background.
@@
    where $p_{1}=(E_{1},|\vec{p_{1}}|)$ and $p_{2}=(E_{2},|\vec{p_{2}}|)$ are the four momentum associated with the two sub-jets (jets) in the final state comprising one (two) central-jet. In events with exactly one central-jet we combine the four momentum of the two sub-jets inside a boosted $\rm{CA}_{1.2}$ jet to get the invariant mass distribution. Whereas in the Nc-jet = 2, case we combine the four momentum of two AK4 jets in order to get the invariant mass associated with the di-jet system. The corresponding distributions in these two final states are plotted in Fig. 13 and 14. One can see that for signal events this distribution has a peak near the Higgs mass. Whereas the same for the background components peak at varying locations, namely, at the mass of the Higgs and $Z$ boson as expected. One also sees a peak at a very low invariant mass which has origin in the final state radiation and hadronization effect from one of the bottom quarks only, and may give rise to jet substructure. In events with two central-jets in final state one has two clear peaks at the values of $Z$ and Higgs mass in the invariant mass distribution corresponding to the background subprocesses. This is due to the presence of $ZZ,\penalty 10000\ Z\nu\bar{\nu}$ component as well as $hZ$ and $h\nu\bar{\nu}$ components of the total SM background. The tail in all these invariant mass distributions at values lower than the peak values is due to the missing momentum component associated with the invisible neutrinos produced in $B$ decays.

### v2

@@ Abstract
Furthermore, this process can also be used to constrain the dark photon mass ($m_{\gamma_{d}}$) and/or the dark gauge coupling ($g_{d}$) consistent with the current muon g2g-2 measurements within the pVLL framework. We perform a detailed collider analysis of the $h\gamma_{d}$ process in the $b\bar{b}\penalty 10000\ +$ missing energy final state. A $2\sigma$ exclusion limit for $m_{\gamma_{d}}$ up to $80$ GeV is obtained assuming $g_{d}=0.05$, $\sin\theta_{L}=4\times 10^{-5}$, $\sin\theta_{s}=0.05$, for a heavy lepton mass $\penalty 10000\ \sim 3$ TeV at a $3$ TeV muon collider with an integrated luminosity of $1$ ab-1.
@@ 1 Introduction
In recent years, the idea of VLLs as portal matter [42, 43, 44, 18] or portal VLLs (pVLLs) has gained significant interest owing to their rich phenomenology, their ability to evade existing collider constraints, and their potential connection to the dark sector, particularly dark matter phenomenology [45, 46].
@@
The main highlight of this work is that the $h\gamma_{d}$ process not only yields a higher Higgs production rate than $hZ$ production at a muon collider, but also offers a noble way to explore dark photons in the mass range 1010100100 GeV, independent of the gauge kinetic mixing parameter ($\varepsilon$) [51, 52, 53, 54]. This is because of the fact that the muon–heavy muon–Higgs vertex is sensitive to the dark photon mass ($m_{\gamma_{d}}$) and the dark gauge coupling ($g_{d}$). Consequently, the absence of any observable enhancement in Higgs production along with missing energy at a muon collider would imply a stringent limit on $m_{\gamma_{d}}$ or $g_{d}$, for fixed values of other relevant model parameters. We illustrate this using several representative benchmark points and show that, even for fermion mixing angle as small as $\sim 10^{-6}$ and heavy muon masses in the range 1133 TeV, a wide range of dark photon masses can be excluded. We illustrarte that a substantial region of the parameter space consistent with current measurement of muon g2g-2 can be probed at the future muon collider using our analysis.
@@ 2 Theoretical Framework
We consider a framework in which the Standard Mode is extended by a vector-like lepton, singlet under $SU(3)_{C}\times SU(2)_{L}$ with $U(1)_{Y}$ hypercharge -$1$. The VLL plays the role of a portal matter, linking the visible SM sector with a dark sector. A simplified realization of such a scenario is to have a VLL transforming non-trivially under an additional local $U(1)_{D}$ symmetry of the dark sector with corresponding gauge boson referred to as the dark photon ($\gamma_{d}$). The dark photon acquires mass through spontaneous breaking of $U(1)_{D}$, achieved by introducing a complex scalar field $\Phi_{d}$ which is a singlet under the SM gauge group but charged under $U(1)_{D}$. Furthermore, if $\Phi_{d}$ carries the same $U(1)_{D}$ charge as the pVLL, a Yukawa-type portal interaction $\omega_{f}\,\Phi_{d}\,\overline{\mu}^{\prime}_{p_{L}}\mu^{\prime}_{R}$ is permitted. The Yukawa interaction reduces to a bilinear term between the pVLL and the SM lepton once $\Phi_{d}$ gets a non-zero VEV. In this work, we focus on the case where the pVLL mixes with the second-generation charged lepton only to avoid constraints coming from flavor violating observables. The assignment of charge of the relevant field content is presented in Table 1.
@@ 2.3 Electroweak precision constraints
The presence of VLL has various observable consequences and can potentially modify the predictions of the SM. It can contribute to the modification of muon life-time, the branching ratio of $Z\to\mu^{+}\mu^{-}$, total width of the Z and W bosons and correction to the oblique parameters etc. Since the VLL carries electromagnetic charge it can modify the Higgs to di-photon decay rate as well. In our model, the pVLL can couple to the Higgs via muon-heavy muon mixing and/or Higgs-dark Higgs mixing. Therefore, the corresponding interaction vertices will be proportional to $\sim(y_{m}\cos\theta_{R}\cos\theta_{L}\cos\theta_{s}\penalty 10000\ +\penalty 10000\ \omega_{f}\cos\theta_{R}\sin\theta_{L}\allowbreak\sin\theta_{s}$). Hence, for small $\sin\theta_{L}$ and $\sin\theta_{s}$ the deviation from the SM $h-\mu-\mu$ vertex factor is less significant. The correction to the oblique parameters, in particular, S and T due to the singlet VLL is expressed as [73]
@@
2.4 Muon g2g-2
The presence of muon–heavy muon–dark photon vertex can give leading contribution to the anomalous muon g2g-2 at one-loop in this pVLL framework. In the limit $m_{\mu_{p}}\approx m_{\mu^{\prime}_{p}}$ and $m_{\mu}\approx m_{\mu^{\prime}}=\dfrac{1}{\sqrt{2}}y_{m}v_{{}_{\rm EW}}$, this contribution is given by [21, 27]
$\Delta a_{\mu}$ $\approx$ $\begin{cases}\penalty 10000\ \penalty 10000\ \dfrac{g^{2}_{d}\,m_{\mu^{\prime}_{p}}\,m_{\mu}}{16\pi^{2}\,m^{2}_{\gamma_{d}}}\sin 2\theta_{R}\,\sin 2\theta_{L},&\penalty 10000\ m_{\mu^{\prime}_{p}}>>m_{\gamma_{d}}\\ \penalty 10000\ \penalty 10000\ \dfrac{g^{2}_{d}\,m_{\mu^{\prime}_{p}}\,m_{\mu}}{4\pi^{2}\,m^{2}_{\gamma_{d}}}\sin 2\theta_{R}\,\sin 2\theta_{L},&\penalty 10000\ m_{\mu}<<m_{\mu^{\prime}_{p}}<<m_{\gamma_{d}}\end{cases}$ (21)
The current world average value of $\Delta a_{\mu}=a^{\rm exp}_{\mu}-a^{\rm SM}_{\mu}$, combining Fermilab Run 1–6 data [75] with the results of Brookhaven experiment [76] is
[…]
which corresponds to $0.6\sigma$ deviation compared to the SM prediction ($a^{\rm SM}_{\mu}=0.00116592033(62)$ [77]).
In Fig. 1, we show the muon g2g-2 allowed region of parameter space both in the $m_{\gamma_{d}}-g_{d}$ and $m_{\mu_{p}}-m_{\gamma_{d}}$ planes consistent with $\Delta a_{\mu}$ mentioned in Eq. (21) at $1\sigma$ level. Fig LABEL:subfig:muon_gm2_mzp_vs_gd corresponds to $\sin\theta_{L}=4\times 10^{-5}$ and four different heavy muon masses (1 TeV, 1.5 TeV, 2 TeV and 3 TeV), whereas Fig LABEL:subfig:muon_gm2_constraint_mzp_vs_mmup corresponds to two distinct sets of $\left(g_{d},\penalty 10000\ \sin\theta_{L}\right)$ with values $\left(0.05,\penalty 10000\ 4\times 10^{-5}\right)$ and $\left(0.5,\penalty 10000\ 1.66\times 10^{-6}\right)$.
Figure 1: Region (coloured) of parameter space allowed by muon g2g-2 measurement at $1\sigma$ level [77] in the (a) $m_{\gamma_{d}}-g_{d}$ plane for $\sin\theta_{L}=4\times 10^{-5}$ and four different $m_{\mu_{p}}$ values (11 TeV, $1.5$ TeV, 22 TeV and 33 TeV) and (b) $m_{\mu_{p}}-m_{\gamma_{d}}$ plane for two sets of ($g_{d},\penalty 10000\ \sin\theta_{L}$).
@@ 3.1 $hZ$ vs $h\gamma_{d}$ at muon collider
Figure 3: Ratios of cross sections ( $R_{i}$’s defined in Eq. (23) ) for different benchmark points at (a) $\sqrt{s}=3$ TeV and (b) $10$ TeV.
One can see that while the cross section for $\mu^{+}\mu^{-}\to hZ$ production in this scenario remains the same as that in the SM, a higher Higgs production rate is predicted in the $\mu^{+}\mu^{-}\to h\gamma_{d}$ channel. This is true even for a $\sin\theta_{L}$ as small [^4] Small $\sin{\theta_{L}}$ ($\sim 10^{-4}$ or less) is favoured by perturbative unitarity requirements for heavy muon mass beyond TeV scale as $\sim 10^{-5}$ with reasonable choices of other relevant parameters (see Table 2 for specification of benchmark points). The underlying reasons behind this is that $\mu_{R}^{\prime}$ and $\mu_{p_{{}_{R}}}^{\prime}$ share identical $SU(2)_{L}\times U(1)_{Y}$ charges. Hence, after transforming them into mass eigen-basis using Eq. (10) no off-diagonal $\mu^{-}_{R}-\mu^{-}_{p_{{}_{R}}}-Z$ interaction arises. However, the absence of $U(1)_{D}$ charge for the muon induces an off-diagonal $\mu_{R}^{-}-\mu^{-}_{p_{{}_{R}}}-\gamma_{d}$ interaction in the mass basis. It turns out that the corresponding interaction vertex $\mu^{+}_{L}-\mu_{p_{{}_{R}}}^{-}-Z$ is not present in this model while the $\mu^{+}_{L}-\mu_{p_{{}_{R}}}^{-}-\gamma_{d}$ (or $\mu^{-}_{R}-\mu_{p_{{}_{L}}}^{+}-\gamma_{d}$ ) vertex factor is proportional to $\sim g_{d}\sin\theta_{R}\cos\theta_{R}$ ( See Eq. (A.2) ). The mixing angle ($\sin\theta_{R}$) between the right chiral muon and heavy muon is given by
@@
In Fig. 4 and Fig. 5, we plot $R_{2}$ which represents the ratio of average unpolarized cross section for $\mu^{+}\mu^{-}\to h\gamma_{d}$ to that of SM $\mu^{+}\mu^{-}\to hZ$ ( defined in Eq. (23) ) in the $m_{\gamma_{d}}$$m_{\mu_{p}}$ plane at $\sqrt{s}=3$ TeV and $\sqrt{s}=10$ TeV, respectively, to illustrate the effect of these two parameters on $R_{2}$ and hence the $h\gamma_{d}$ cross section for two different $g_{d}$ and $\sin\theta_{L}$ and fixed $\sin\theta_{s}$. The perturbative unitarity limit obtained using Eq. (15) is also displayed on the same plot. Since the requirement of perturbative unitarity disfavors low values of $v_{d}$, one cannot achieve an arbitrary high values of $R_{2}$ by going to lower values of dark photon mass or higher values of $g_{d}$ and a trade off is maintained in the choice of the two independent parameters $m_{\gamma_{d}}$ and $g_{d}$}.
Figure 4: Variation of ratio of cross sections ($R_{2}$) for the process $\mu^{+}\mu^{-}\to h\gamma_{d}$ with respect to the SM $hZ$ production rate in $m_{\gamma_{d}}$$m_{\mu_{{}_{p}}}$ plane assuming $\sqrt{s}=3$ TeV for two different set of ($g_{d},\penalty 10000\ \sin\theta_{L}$) and a fixed value of $\sin\theta_{s}$. (a) The left plot corresponds to $g_{d}=0.05,\penalty 10000\ \sin\theta_{s}=0.05,\penalty 10000\ \text{and}\penalty 10000\ \sin\theta_{L}=4\times 10^{-5}$ and (b) the right plot corresponds to $g_{d}=0.5,\penalty 10000\ \sin\theta_{s}=0.05,\penalty 10000\ \text{and}\penalty 10000\ \sin\theta_{L}=1.66\times 10^{-6}$. The region above the dashed (black) line is allowed by the constraint on $\omega_{f}$ described in Eq. (15).
Figure 5: Variation of ratio of cross sections ($R_{2}$) for the process $\mu^{+}\mu^{-}\to h\gamma_{d}$ with respect to the SM $hZ$ production rate in $m_{\gamma_{d}}$$m_{\mu_{{}_{p}}}$ plane assuming $\sqrt{s}=10$ TeV for two different set of ($g_{d},\penalty 10000\ \sin\theta_{L}$) and a fixed value of $\sin\theta_{s}$. (a) The left plot corresponds to $g_{d}=0.05,\penalty 10000\ \sin\theta_{s}=0.05,\penalty 10000\ \text{and}\penalty 10000\ \sin\theta_{L}=4\times 10^{-5}$ and (b) the right plot corresponds to $g_{d}=0.5,\penalty 10000\ \sin\theta_{s}=0.05,\penalty 10000\ \text{and}\penalty 10000\ \sin\theta_{L}=1.66\times 10^{-6}$. The region above the dashed (black) line is allowed by the constraint on $\omega_{f}$ described in Eq. (15).
@@ 3.2 Analysis of $h\gamma_{d}$ process
The cross section for $h\gamma_{d}$ production strongly depends on dark photon mass and dark gauge coupling as well as on other model parameters (see Fig. 4, Fig. 5, as well as the corresponding vertex factors in Appendix A and the squared matrix elements in Appendix B). The cross section for the $\mu^{+}\mu^{-}\to h\gamma_{d}$ process as a function of the dark gauge coupling ($g_{d}$) can be parametrized as
@@
As already mentioned, Higgs boson produced at the future muon collider center-of-mass energies in a $2\to 2$ process are typically highly boosted. The Higgs subsequently decays into a pair of bottom quarks which due to their small angular separation, manifest as a single collimated jet in the detector. In Fig. 11 and 12, we plot the transverse momentum distributions of a jet or di-jet system, as applicable, for two different muon collider center-of-mass energies. The purpose of these distributions are two-fold. They provide an explanation of Fig. 9 and 10 and also give a motivation for the following discussion.
@@ 3.3 Kinematic variables
    In Fig. 13, we plot the missing energy distributions for the signal events for various benchmark scenarios corresponding to two different final states at center-of-mass energy $\sqrt{s}=3\penalty 10000\ \text{TeV}$. The corresponding distributions at 10 TeV center-of-mass energy are shown in Fig. 14. The contribution to the missing energy for the signal events is due to the invisible dark photon and the distribution is expected to peak $\sim{\sqrt{\rm s}/2}$ (neglecting the finite mass effect) in a $2\to 2$ process. However, due to the presence of additional neutrinos coming from $B$ meson decays the missing energy distribution extends beyond $\sqrt{\rm s}/2$}. The SM background events in the $b\bar{b}\nu\bar{\nu}$ final state has dominant contribution from the $2\to 3$ processes such as $h\nu\bar{\nu}$ and $Z\nu\bar{\nu}$. Together they form the 99 % of the total cross section and only a very small fraction of it comes from $2\to 2$ processes like $ZZ$ and $hZ$. The pair of neutrinos in the $h\nu\bar{\nu}$ or $Z\nu\bar{\nu}$ processes carries maximum share of the available center-of-mass energy. Hence, the missing energy distribution for the SM background events has a peak mostly around $\sqrt{s}$. This is reflected in the missing energy distributions for individual SM background components and the total SM background as depicted in the same figures. Therefore, missing energy variable is an extremely useful discriminator between the signal and the background.
@@
    where $p_{1}=(E_{1},|\vec{p_{1}}|)$ and $p_{2}=(E_{2},|\vec{p_{2}}|)$ are the four momentum associated with the two sub-jets (jets) in the final state comprising one (two) central-jet. In events with exactly one central-jet we combine the four momentum of the two sub-jets inside a boosted $\rm{CA}_{1.2}$ jet to get the invariant mass distribution. Whereas in the Nc-jet = 2, case we combine the four momentum of two AK4 jets in order to get the invariant mass associated with the di-jet system. The corresponding distributions in these two final states are plotted in Fig. 15 and 16. One can see that for signal events this distribution has a peak near the Higgs mass. Whereas the same for the background components peak at varying locations, namely, at the mass of the Higgs and $Z$ boson as expected. One also sees a peak at a very low invariant mass which has origin in the final state radiation and hadronization effect from one of the bottom quarks only, and may give rise to jet substructure. In events with two central-jets in final state one has two clear peaks at the values of $Z$ and Higgs mass in the invariant mass distribution corresponding to the background subprocesses. This is due to the presence of $ZZ,\penalty 10000\ Z\nu\bar{\nu}$ component as well as $hZ$ and $h\nu\bar{\nu}$ components of the total SM background. The tail in all these invariant mass distributions at values lower than the peak values is due to the missing momentum component associated with the invisible neutrinos produced in $B$ decays.